# EMNLP 2024 Paraphrase Types Elicit Prompt Engineering Capabilities

### v4

@@ Appendix C Additional Results
Prompt Complexity Correlations. Table 2 shows the individual correlations between the three markers of prompt complexity and downstream task performance as described in Section B.2. All correlations are close to zero, and only very small correlations exist for question answering (0.13 - 0.14) and sentence composition (0.10 - 0.11) with $p<0.05$. We could generally find no evidence that the complexity of the prompt contributes to gains or losses in performance outcomes.
Prompt Lengths. We conducted a post-hoc analysis of the number of input tokens across various task families using the GPT-2 tokenizer. As shown in Table 3, the average number of tokens varies significantly depending on the nature of the task, ranging from 36 tokens for title generation tasks to 390 tokens for text-to-code tasks. The overall distribution of input tokens is even broader when looking at individual tasks, with the shortest task, sentence composition, using as few as 19 tokens, and the longest, question rewriting, reaching up to 753 tokens. This wide spectrum of prompt lengths indicates that our study covers a diverse array of real-world use cases, from concise to extended prompts. By analyzing 120 tasks across 24 task families, we believe our conclusions hold across a broad range of prompt lengths, including those much shorter (e.g., fewer than 36 tokens) and much longer (e.g., more than 800 tokens).
Results by Task and Type. We show a detailed table of avgerage downstream task performance gain or loss over the original prompt of a task when paraphrasing the prompt using modifications of one of the six different paraphrase groups in Table 10. We decompose this table for each model in Tables 12, 11, 13, 15 and 14.
Maximum Potential Gain. To understand how much models would possibly benefit from adjusting prompts in optimal ways as explained in Table 8, we also show the maximum potential gain in Figure 9 of the teaser on the first page of this paper (Figure 1). This gives an upper bound to what is possible if prompts were optimally adjusted with paraphrase types. Again, we note that although finding these optimal prompt adjustments is challenging, this task is not impossible and gives space for future work.
Model Scale Comparison. Figure 12 compares model size between LLaMA 3 8B and LLaMA 3 70B. Observe how the smaller 8B parameter model benefits more across tasks when compared to its larger 70B version. Only for program execution, information extraction, sentence completion, question rewriting, and question generation, the 70B model shows a larger gain.
@@
Figure 11: The average percentage in downstream task performance gain or loss for applying different paraphrase types across all five models. This decomposes the groups of Figure 3 into their individual types.
@@
Table 8: How we calculate the potential performance gain for paraphrasing prompts using a selected example from the Amazon food review summarization task McAuley and Leskovec (2013) with LLaMA 3 8B. Median Gain. To calculate the median gain (as in Figure 1), we consider all paraphrases of the prompt that are better than the original prompt (green and yellow). Median Loss. To calculate the median loss (as in Figure 1), we consider all paraphrases of the prompt that are worse than the original prompt (red). Max. We calculate the maximum possible gain by aggregating only the best performing paraphrases of the prompts highlighted in yellow.

### v5

@@ Appendix C Additional Results
Prompt Complexity Correlations. Table˜2 shows the individual correlations between the three markers of prompt complexity and downstream task performance as described in Section˜B.2. All correlations are close to zero, and only very small correlations exist for question answering (0.13 - 0.14) and sentence composition (0.10 - 0.11) with $p<0.05$. We could generally find no evidence that the complexity of the prompt contributes to gains or losses in performance outcomes.
Prompt Lengths. We conducted a post-hoc analysis of the number of input tokens across various task families using the GPT-2 tokenizer. As shown in Table˜3, the average number of tokens varies significantly depending on the nature of the task, ranging from 36 tokens for title generation tasks to 390 tokens for text-to-code tasks. The overall distribution of input tokens is even broader when looking at individual tasks, with the shortest task, sentence composition, using as few as 19 tokens, and the longest, question rewriting, reaching up to 753 tokens. This wide spectrum of prompt lengths indicates that our study covers a diverse array of real-world use cases, from concise to extended prompts. By analyzing 120 tasks across 24 task families, we believe our conclusions hold across a broad range of prompt lengths, including those much shorter (e.g., fewer than 36 tokens) and much longer (e.g., more than 800 tokens).
Results by Task and Type. We show a detailed table of avgerage downstream task performance gain or loss over the original prompt of a task when paraphrasing the prompt using modifications of one of the six different paraphrase groups in Table˜10. We decompose this table for each model in Tables˜12, 11, 13, 15 and 14.
Maximum Potential Gain. To understand how much models would possibly benefit from adjusting prompts in optimal ways as explained in Table˜8, we also show the maximum potential gain in Figure˜9 of the teaser on the first page of this paper (Figure˜1). This gives an upper bound to what is possible if prompts were optimally adjusted with paraphrase types. Again, we note that although finding these optimal prompt adjustments is challenging, this task is not impossible and gives space for future work.
Model Scale Comparison. Figure˜12 compares model size between LLaMA 3 8B and LLaMA 3 70B. Observe how the smaller 8B parameter model benefits more across tasks when compared to its larger 70B version. Only for program execution, information extraction, sentence completion, question rewriting, and question generation, the 70B model shows a larger gain.
@@
Figure 11: The average percentage in downstream task performance gain or loss for applying different paraphrase types across all five models. This decomposes the groups of Figure˜3 into their individual types.
@@
Table 8: How we calculate the potential performance gain for paraphrasing prompts using a selected example from the Amazon food review summarization task McAuley and Leskovec (2013) with LLaMA 3 8B. Median Gain. To calculate the median gain (as in Figure˜1), we consider all paraphrases of the prompt that are better than the original prompt (green and yellow). Median Loss. To calculate the median loss (as in Figure˜1), we consider all paraphrases of the prompt that are worse than the original prompt (red). Max. We calculate the maximum possible gain by aggregating only the best performing paraphrases of the prompts highlighted in yellow.